At the school carnival, Jade has a game booth with a spinner. The banner at her booth reads "20% chance of winning!" What is the chance that a player may not win?
They have an 80% chance of loosing. Just subtract 20 from 100 and you are left with 80.
100% - 20% = 80%
There is 100% chance FULLY so 100 - 20 = 80. 80% chance.